### v1

@@ 2.2 Bayesian Optimal Experimental Design
3:Train GP using the dataset $D=\{\mathbf{X}_{\text{init}},\mathbf{Y}_{\text{init}}\}$.
4:for $i=1$ to $N-n_{0}$ do $\triangleright$ Remaining simulation budget
5:  Initialize a candidate input $\mathbf{x}_{\text{cand}}$ within $\Xi$.
6:  for iteration = 1 to desired optimization steps do $\triangleright$ Optimize EIG
7:   Compute the Expected Information Gain (EIG) for $\mathbf{x}_{\text{cand}}$ using Eq. (8).
8:   Update $\mathbf{x}_{\text{cand}}$ to improve the EIG.
9:  end for
10:  Select the optimized input $\mathbf{x}^{*}$ with the highest EIG.
11:  Run the simulation $\mathcal{G}$ at $\mathbf{x}^{*}$ to obtain its output $y^{*}$.
12:  Add $\{\mathbf{x}^{*},y^{*}\}$ to the dataset $D$.
13:  Refit GP using the updated dataset $D$.
14:end for
15:Solve the Bayesian calibration problem using MCMC with the GP emulator.
Algorithm 2 Goal-Oriented Bayesian Optimal Experimental Design (GBOED)
Input: Bounds for input space $\Xi$, initial number of samples $n_{0}$, simulation budget $N$, GP emulator GP
@@
3:Train GP using the dataset $D=\{\mathbf{X}_{\text{init}},\mathbf{Y}_{\text{init}}\}$.
4:for $i=1$ to $N-n_{0}$ do $\triangleright$ Remaining simulation budget
5:  Solve the calibration problem using MCMC with the current surrogate model:
$\mathbf{X}^{\prime}\sim p(\theta\mid y_{\text{obs}},\mathcal{GP}(D))$
6:  Initialize a candidate input $\mathbf{x}_{\text{cand}}$ within $\Xi$.
7:  for iteration = 1 to desired optimization steps do $\triangleright$ Optimize EIG
8:   Compute the Expected Information Gain (EIG) for $\mathbf{x}_{\text{cand}}$ with respect to the predictions at $\mathbf{X}^{\prime}$ via Eq. (8).
9:   Update $\mathbf{x}_{\text{cand}}$ to maximize $EIG(\mathbf{x}_{\text{cand}})$.
10:  end for
11:  Select the optimized input $\mathbf{x}^{*}$ with the highest EIG.
12:  Run the simulation $\mathcal{G}$ at $\mathbf{x}^{*}$ to obtain its output $y^{*}$.
13:  Add $\{\mathbf{x}^{*},y^{*}\}$ to the dataset $D$.
14:  Refit GP using the updated dataset $D$.
15:end for
16:Solve the final Bayesian calibration problem using MCMC using the GP emulator to obtain posterior samples:
@@ 2.3 Calibrate, emulate, sample
At a high level, CES starts with a calibration phase, in which ensemble Kalman filter techniques such as Ensemble Kalkman Inversion (EKI) [28] or Ensemble Kalman Sampling (EKS) [19], are used to obtain an initial posterior distribution of parameters based on a limited number of model evaluations. The next phase, Emulate, involves training an efficient GP emulator on parameter samples (inputs) and model outputs(s) obtained during the Calibration phase. This emulator is a computationally cheap surrogate for $\mathcal{G}$.
@@
We implement CES following methods specified in [7]. We use EKS for the calibration phase, since EKS (unlike EKI) does not suffer from the collapse of the posterior distribution to a point mass [19]. EKS is used only for a limited number of iterations to reduce the computational burden, so the subsequent steps of CES (Emulation and Sampling) are usually necessary to obtain reliable UQ and refinement of the posterior distribution of $\theta$.
@@ 2.4 History Matching
HM is a statistical methodology that leverages observed data to refine model parameters by ruling out “implausible” parameter values based on how well they match the observations [56]. By iteratively updating parameter configurations, HM effectively narrows the range of plausible values, thereby reducing the computational costs associated with the expensive model simulations. History matching is well-represented in the field of climate science and has been used extensively for climate model calibration [8, 57, 56, 16, 47, 31].
@@
Subsequent waves refine the NROY space further. For each wave, new training samples are drawn from within the current NROY space, the expensive model is run, and the emulator is updated. Sampling from the NROY space is challenging because uniform sampling is infeasible when the space becomes fragmented due to eliminated “clusters of inputs”. Instead, we sample uniformly from the remaining points already identified as part of the NROY space. The HM process continues until a stopping criterion is met. This could be based on computational constraints or when the NROY space becomes too small to warrant further refinement. In our implementation, we use a fixed threshold of three for the implausibility metric in all waves and rely on the available computational budget to determine when to stop. To automate the HM process, [10] proposes a semi-automated approach using sequential MCMC, which eliminates the need for user-specified parameters like implausibility thresholds. Instead, we adhere to a simpler framework with fixed thresholds and computational constraints guiding our stopping criteria. Importantly, while the HM algorithm technically ends when the stopping criteria is reached, we apply an extra MCMC step, based on the GP emulator fitted to samples in the final wave. Unlike traditional HM, the final MCMC step allows us to construct a posterior distribution for calibration parameters.
@@
3 Experimental setup
We use the Lorenz ’96 multiscale system described in [35] (Eq. 10) to test our methods. The Lorenz ’96 system is a simplified mathematical model that is often used to study the behavior of chaotic systems such as climate [32, 41, 40, 43]. The system consists of a set of ordinary differential equations that describe the evolution of a large number of interacting variables. In this system, there are $K$ large-scale variables ($X_{k}$) and $JK$ small-scale variables ($Y_{j,k}$). The $X$ and $Y$ variables interact with each other, but there is also coupling between them, as shown in Eq. (10).
@@ 3.2 Implementation details
To assess how the performance of methods changed with increasing model evaluations, we performed calibration twice with each method, using relatively few (200) and more (1000-5400) model evaluations. We chose 200 model evaluations based on methodology in [59], and 1000 represents a reasonable upper bound for the number of simulations that would be practically feasible with an expensive climate model [11, 13]. To obtain the final posterior distribution for $\theta$ for all methods, we use GP emulators, along with corresponding $\phi_{k}(Z)$ and $\Gamma_{obs}$ to perform MCMC sampling. Recall that because the observables are five-dimensional, five independent GPs were fitted. We use the same prior specification for GP lengthscales as specified in [7] for all methods for consistency. MCMC sampling was conducted via Sequential Tempered Markov Chain Monte Carlo (ST-MCMC) [4], which is a highly effective Sequential Monte Carlo (SMC) sampler due to its efficient parallelism while retaining sampling robustness due to its adaptivity and monitoring of sample population statistics instead of individual chain statistics. We found that 8192 ST-MCMC parallel samples resulted in a comprehensive exploration of the posterior for all methods. See [4] for more details on ST-MCMC sampling.
@@ 3.2.1 CES and EKS
For performing Bayesian inversion on the Lorenz ’96 system (Eq. 10) via CES, EKS was run for 54 iterations with 100 ensemble members (5400 model evaluations), in accordance with [7]. We also ran EKS with only 10 iterations and 20 ensemble members (200 model evaluations). For the case of 5400 model evaluations, only iterations $1,6,12,\dots,54$ were saved for subsequent GP fitting, as suggested in [7]. Doing so reduces the total number of data required to fit GPs, decreasing computation time while retaining the spread of the samples over the total number of EKS iterations.
@@ 3.2.2 BOED and GBOED
For both BOED and GBOED with 200 or 1000 total evaluations, the initial GP model was trained using 64 Quasi Monte Carlo samples from the Sobol Sequence to ensure space-filling coverage of the input domain. These initial samples guided subsequent greedy OED choices. The remaining samples were selected using the optimization procedures outlined in Algorithms 2.1 and 2.2, with ST-MCMC employed for the intermediate sampling required by GBOED.
@@ 4 Experiments
In this section, we describe three experiments pertaining to calibration of the Lorenz ’96 multiscale system using the main methods CES, HM, BOED, and GBOED. We begin with examining the marginal posterior distributions of $h,F$, $c$ and $b$. Next, we conduct a convergence analysis of all methods over increasing model evaluations, and finally, we conduct an in-depth performance assessment comparing accuracy and UQ of all methods.
@@
First, we show the posterior distribution resulting from all methods for $\theta=(1,10,10,10)$. Posterior distributions for $h,F,c$ and $b$ for the case of fewer (200) and more (1000-5400) model evaluations resulting from the main methods CES, HM, BOED, and GBOED are shown in the left and right panels, respectively, of Figure 1 (recall that CES was run with 5400 model evaluations following methods in [7]). Due to the nonlinear and chaotic nature of the Lorenz ’96 system and correlation among parameters, the resulting posterior distributions are not expected to be Gaussian. Indeed, our MCMC samples often exhibit skewness and heavy tails.
@@
While posterior distributions for each of the four parameters appear similar across methods, there are some differences. Notably, the ranges for parameter $c$ for CES, BOED, and GBOED are much wider (0-50) compared to that of HM (0-20), and this is consistent regardless of the number of model evaluations used. Posteriors for parameter $b$ tend toward Gaussian, but posteriors for $F$ and $c$ are right-skewed. Posteriors for parameter $h$ appear somewhat uniform (HM1000, GBOED1000), slightly bimodal (GBOED200, BOED1000), or approximately normal (CES200, CES5400). Despite differences in posteriors, all methods were able to capture the true parameter values (denoted by black vertical lines in Figure 1). Note also that as the number of model evaluations increases, it may not necessarily be the case that the variance of the posterior distribution will decrease. As model evaluations increase, methods such as CES and GBOED especially may explore previously under-sampled regions of parameter space, leading to a more realistic representation of uncertainty.
@@
Figure 1: Frequency histograms of posterior distributions of parameters $h,F,c$, and $b$ resulting from CES200, HM200, BOE200, and GBOED200 (left) and CES1000, HM1000, BOE1000, and GBOED1000 (right). Histograms represent 8192 values. True parameter values are denoted by vertical lines.
4.1 Convergence analysis
@@
Next we conducted an exercise to gain insight into convergence rates of the main methods in calibrating the Lorenz ’96 system. The goal of the convergence exercise was to determine how performance metrics of each method changed over increasing model evaluations. We report both score and AE every 100 evaluations up to 1000 total model evaluations for all methods. We included the intermediate (calibration) step of CES (EKS) as well for comparison.
AEs generally decreased as the number of model evaluations increased for parameter $b$, but there was considerable variability among methods for parameters $h$, $F$ and $c$ (Figure 2, left panel). For instance, AEs for parameter $c$ declined for methods BOED, CES, and EKS, but AEs increased markedly after 350 model evaluations for GBOED and were nearly constant for HM. In fact, AEs for GBOED consistently increased over model evaluations for all parameters, suggesting potential overfitting or instability.
For scores, the behavior varied by method. CES, EKS, BOED, and to some extent HM demonstrated slight but consistent improvements in score over increasing model evaluations (Figure 2, right panel). The stability of scores for HM over increasing model evaluations suggests that HM is less sensitive to the quantity of model evaluations once a sufficient baseline of information has been established. GBOED was unique in exhibiting a gradual decline in scores (degradation in performance) for all parameters over the 1000 model evaluations, which may be due to optimization quality and hyperparameter sensitivity (see SM3).
@@
Figure 2: AEs (left) and scores (right) for CES, EKS, HM, BOED, and GBOED over increasing model evaluations for calibrating the Lorenz ’96 system using base parameters $\theta=(1,10,10,10)$.
This convergence analysis suggests that CES, GBOED, and HM may require as few as 500 model evaluations to achieve accurate parameter estimates with robust uncertainty quantification. However, the results also highlight important differences in behavior across methods. For example, while BOED and EKS appear to benefit from additional evaluations (with metrics continuing to improve at 1000 evaluations), GBOED performance may be hindered by excessive model evaluations.
Finally, it is important to note that these conclusions are based on a single set of true parameters for the Lorenz ’96 system. The convergence behavior might differ for other parameter configurations, and future work should explore this sensitivity to provide a more comprehensive understanding of method performance across a variety of scenarios.
4.2 Performance evaluation
@@
To obtain a fair and rigorous comparison of the main methods, we performed Bayesian inversion for $\theta_{i},i=1\dots 60$ using the Lorenz ’96 multiscale system (Eq. 10). Each component of $\theta_{i}$ was drawn from the following prior distributions based on baseline parameter values of 1, 10, log10, and 10 for $h,F$, $c$, and $b$, respectively:
@@ 4.2.1 Statistical analyses
where, in this formulation, $\beta_{1}$ represents the effects of the 10 levels of $METHOD$ relative to the referencereference level, and $\beta_{2}$ represents the effects of the three levels of $PARAM$ relative to the reference level. This model can be generally expressed in matrix form as $\mathbf{Y}=\mathbf{X}\boldsymbol{\beta}+\boldsymbol{\epsilon}$, where $\mathbf{X}$ is a design matrix that encodes the levels of $METHOD$ and $PARAM$. The coefficient vector $\boldsymbol{\beta}$ contains ten coefficients for $METHOD$, three coefficients for $PARAM$, and one for the intercept. Additionally, we excluded extremely low score values (less than -50, primarily from LHS methods) from the dataset, as these could have disproportionately influenced the results. In total, 20 observations were removed. After fitting these models, we performed post-hoc pairwise comparisons of the estimated marginal means of AE by $METHOD$ to determine if differences between any two calibration methods were significantly different. We applied the Bonferroni correction for post-hoc pairwise comparisons to control for family-wise error rate (FWER). While Tukey’s Honest Significant Difference (HSD) is commonly used for ANOVA post-hoc tests, it assumes equal group sizes, whereas Bonferroni remains valid for unequal sample sizes. Given our study design, Bonferroni provides a more conservative assessment of significant differences, which is what we desire. Significance of pairwise tests was assessed at $\alpha=0.05$. All models were fitted in R using the lm() function [45]; pairwise comparisons were carried out using the emmeans [29] and modelbased [37] packages.
4.2.2 Results
Overall, EKS5400 and GBOED200 outperformed all other methods with respect to AE and score for calibration of the Lorenz ’96 system. Most methods successfully estimated parameter $h$ with a higher accuracy compared to parameters $F,c$ and $b$ (Figure 3 A). The largest variability in performance was exhibited in parameter $c$, with AEs reaching values as high as 20 for some methods (Figure 3 A). LHS-based methods consistently underperformed in terms of both accuracy and UQ, emphasizing the limitations of relying solely on prior-based sampling without incorporating a calibration step (Figures 3 A and B).
Figure 3: Boxplots of AEs (A) and log score (B) for calibrated parameters $h,F,c$ and $b$ of the Lorenz ’96 model for all method. Each boxplot for AE represents 60 values; values less than -50 for score were omitted. AE values were calculated using the mean of the posterior distribution. Horizontal lines within boxplots represent the median, while stars denote means. Parameter $c$ has been back-transformed from a log scale.
Pairwise comparisons of estimated marginal means of AE (averaged over all parameters) from the linear AE model (Eq. 12) revealed notable performance trends. LHS-based methods (LHS200, LHS1000) and BOED200 performed significantly worse than all other methods (Table 1, Figure 4). EKS5400 and GBOED200 outperformed six other methods including BOED200, BOED1000, CES200, EKS200, LHS200, and LHS1000 (Figure 4, Table 1). CES5400 and GBOED1000 also performed well but did not significantly outperform EKS200 (Table 1).

### v2

@@ 2.2 Bayesian Optimal Experimental Design
3:Train GP using the dataset $D=\{\mathbf{X}_{\text{init}},\mathbf{Y}_{\text{init}}\}$ (including hyperparameter estimation).
4:Let $K=\left\lceil\frac{N-n_{0}}{B}\right\rceil$ denote the number of acquisition batches.
5:for $k=1$ to $K$ do $\triangleright$ Batched acquisitions
6:  Initialize a batch of candidate inputs $\mathbf{X}_{\text{cand}}\subset\Xi$ with $|\mathbf{X}_{\text{cand}}|=B$.
7:  for iteration = 1 to desired optimization steps do $\triangleright$ Optimize batch EIG
8:   Compute the Expected Information Gain (EIG) for $\mathbf{X}_{\text{cand}}$ using Eq. (8).
9:   Update $\mathbf{X}_{\text{cand}}$ to improve the EIG.
10:  end for
11:  Select the optimized batch $\mathbf{X}^{*}$ with the highest EIG.
12:  Run the simulation $\mathcal{G}$ for each input in $\mathbf{X}^{*}$ to obtain outputs $\mathbf{Y}^{*}$.
13:  Add $\{\mathbf{X}^{*},\mathbf{Y}^{*}\}$ to the dataset $D$.
14:  Refit GP hyperparameters using the updated dataset $D$.
15:end for
16:Solve the Bayesian calibration problem using MCMC with the GP emulator.
Algorithm 2 Goal-Oriented Bayesian Optimal Experimental Design (GBOED) with batched acquisitions
Input: Bounds for input space $\Xi$, initial number of samples $n_{0}$, total simulation budget $N$, batch size $B$, GP emulator GP
@@
1:Select an initial set of $n_{0}$ input samples $\mathbf{X}_{\text{init}}\subset\Xi$.
@@
3:Train GP on $D=\{\mathbf{X}_{\text{init}},\mathbf{Y}_{\text{init}}\}$ (including hyperparameter estimation).
4:Let $K=\left\lceil\frac{N-n_{0}}{B}\right\rceil$ denote the number of acquisition batches.
5:for $k=1$ to $K$ do $\triangleright$ Batched acquisitions
6:  Solve the calibration problem using MCMC with the current surrogate model to obtain posterior samples:
$\mathbf{X}^{\prime}\sim p(\theta\mid y_{\text{obs}},\mathcal{GP}(D)).$
7:  Initialize a candidate batch $\mathbf{X}_{\text{cand}}\in\Xi^{B}$ (i.e., $B$ candidate inputs).
8:  for iteration = 1 to desired optimization steps do $\triangleright$ Optimize batch EIG
9:   Compute the goal-oriented expected information gain for the batch $\mathbf{X}_{\text{cand}}$ with respect to predictions at $\mathbf{X}^{\prime}$ via Eq. (8).
10:   Update $\mathbf{X}_{\text{cand}}$ to increase $EIG(\mathbf{X}_{\text{cand}})$.
11:  end for
12:  Select the optimized batch $\mathbf{X}^{*}=\{\mathbf{x}_{1}^{*},\ldots,\mathbf{x}_{B}^{*}\}$.
13:  Run $\mathcal{G}$ at each $\mathbf{x}_{b}^{*}$ to obtain outputs $\mathbf{Y}^{*}=\{y_{1}^{*},\ldots,y_{B}^{*}\}$.
14:  Augment the dataset: $D\leftarrow D\cup\{\mathbf{X}^{*},\mathbf{Y}^{*}\}$.
15:  Refit GP on the updated dataset $D$ (including GP hyperparameter re-estimation).
16:end for
17:Solve the final Bayesian calibration problem using MCMC with the final surrogate:
@@ 2.3 Calibrate, emulate, sample
At a high level, CES starts with a calibration phase, in which ensemble Kalman filter techniques such as Ensemble Kalman Inversion (EKI) [44] or Ensemble Kalman Sampling (EKS) [27], are used to obtain an initial posterior distribution of parameters based on a limited number of model evaluations. The next phase, Emulate, involves training an efficient GP emulator on parameter samples (inputs) and model outputs(s) obtained during the Calibration phase. This emulator is a computationally cheap surrogate for $\mathcal{G}$.
@@
We implement CES following methods specified in [11]. In this study we use EKS and EKI as described in [37] for our examples. EKS (unlike most EKI algorithms) does not suffer from collapse of the posterior distribution, but it is computationally expensive and may require more iterations than EKI [27]. EKS is typically used only for a limited number of iterations to reduce the computational burden, so the subsequent steps of CES (Emulation and Sampling) are necessary to obtain reliable UQ and refinement of the posterior distribution. We note that it is also possible to use alternative non-collapsing or regularized variants of EKI. A number of such variants have been developed that help preserve ensemble spread and offer fast convergence. TEKI (Tikhonov-regularized EKI) adds explicit quadratic regularization to the update, mitigating collapse and stabilizing inference in ill-posed settings [8]. In addition, Kalman–Bucy/continuous-time EKI with early-stopping rules has been analyzed as a data-driven regularization that halts the dynamics before ensemble collapse [67]. These variants yield final ensembles whose spread better reflects posterior uncertainty and improving computational efficiency relative to running many small EKS steps [67]. Finally, statistically-linearized EKI variants (EKI-SL/IEKF-SL) explicitly prevent collapse in linear tests and maintain empirical covariances near posterior levels [7]. Finally, [36] describe an iterated Kalman framework in which algorithmic choices (discretization, step control) can influence ensemble collapse and computational efficiency.
@@ 2.4 History Matching
HM is a statistical methodology that leverages observed data to refine model parameters by ruling out “implausible” parameter values based on how well they match the observations [73]. By iteratively updating parameter configurations, HM effectively narrows the range of plausible values, thereby reducing the computational costs associated with the expensive model evaluations. History matching is well-represented in the field of climate science and has been used extensively for climate model calibration [14, 74, 73, 22, 61, 46].
@@
Subsequent waves refine the NROY space further. For each wave, new training samples are drawn from within the current NROY space, the expensive model is run, and the emulator is updated. Sampling from the NROY space is challenging because uniform sampling is infeasible when the space becomes fragmented due to eliminated “clusters of inputs”. Instead, we sample uniformly from the remaining points already identified as part of the NROY space. The HM process continues until a stopping criterion is met. This could be based on computational constraints or when the NROY space becomes too small to warrant further refinement. In our implementation, we use a fixed threshold of three for the implausibility metric in all waves and rely on the available computational budget to determine when to stop. To automate the HM process, [18] proposes a semi-automated approach using sequential MCMC, which eliminates the need for user-specified parameters like implausibility thresholds. Instead, we adhere to a simpler framework with fixed thresholds and computational constraints guiding our stopping criteria.
The HM algorithm is performed first to generate a set of informative input–output pairs. After the algorithm terminates, all inputs and corresponding model outputs generated across all waves are collected, and Gaussian process emulators are fitted to these data, as described in Section 2.1. The final emulators are then used within an MCMC framework to construct the posterior distribution over calibration parameters. Within this final MCMC step, both the GP predictive mean and predictive covariance are used to define the likelihood (Eq. 5).
@@
3 Lorenz’96 Multiscale system
We first use the Lorenz ’96 multiscale system described in [49] (Eq. 10) to test calibration methods. The Lorenz ’96 multiscale system is often used to study the behavior of chaotic systems such as climate [46, 55, 54, 57]. The system consists of a set of ordinary differential equations that describe the evolution of a large number of interacting variables. In this system, there are $K$ large-scale variables ($X_{k}$) and $JK$ small-scale variables ($Y_{j,k}$). The $X$ and $Y$ variables interact with each other, but there is also coupling between them, as shown in Eq. (10).
@@ 3.2 Implementation details
To assess how the performance of methods changed with increasing model evaluations, we performed calibration twice with each method, using relatively few (200) and more (1000-5400) model evaluations. We chose 200 model evaluations based on methodology in [76], and 1000 represents a reasonable upper bound for the number of model evaluations that would be practically feasible with an expensive climate model [19]. To obtain the final posterior distribution for $\theta$ for all methods, we use GP emulators, along with corresponding $\phi_{k}(Z)$ and $\Gamma_{obs}$ to perform MCMC sampling. We use the same prior specification for GP lengthscales as specified in [11] for all methods for consistency. MCMC sampling was conducted via Sequential Tempered Markov Chain Monte Carlo (ST-MCMC) [6], which is a highly effective Sequential Monte Carlo (SMC) sampler due to its efficient parallelism while retaining sampling robustness due to its adaptivity and monitoring of sample population statistics instead of individual chain statistics. We found that 8092 ST-MCMC parallel samples resulted in a comprehensive exploration of the posterior for all methods. See [6] for more details on ST-MCMC sampling.
@@ 3.2.1 CES and EKS
For both examples, we use EKS with an adaptive timestep as described in [27], which is also the version implemented in the original CES paper [11]. EKS was implemented using the Julia package EnsembleKalmanProcesses.jl [12] using default settings. We did not impose strict bounds on the prior in our implementation of EKS. For performing Bayesian inversion on the Lorenz ’96 system (Eq. 10) via CES, EKS was run for 54 iterations with 100 ensemble members (5400 model evaluations), in accordance with [11]. We also ran EKS with only 10 iterations and 20 ensemble members (200 model evaluations). For the case of 5400 model evaluations, only iterations $1,6,12,\dots,54$ were saved for subsequent GP fitting, also as suggested in [11]. Doing so reduces the total number of data required to fit GPs, decreasing computation time while retaining the spread of the samples over the total number of EKS iterations.
@@ 3.2.2 BOED and GBOED
For both BOED and GBOED with 200 or 1000 total model evaluations, the Gaussian process (GP) surrogate was constructed in a batched sequential design framework. An initial GP model was trained using 20 prior Monte Carlo samples for coverage of the input domain. Thereafter, additional design points were selected in batches of 20, with the GP surrogate retrained after each batch. Within each iteration, candidate design points were optimized using L-BFGS-B with gradients computed using JAX automatic differentiation. These steps were repeated until the target evaluation budget was reached as described in Algorithms 1 and 2. For GBOED, the intermediate sampling required to approximate the expected information gain was performed using ST-MCMC [6].
@@ 3.3 Experiments
In this section, we describe experiments pertaining to calibration of the Lorenz ’96 multiscale system using the main methods CES, HM, BOED, and GBOED. First, we examine the marginal posterior distributions of $h,F$, $c$ and $b$. We then conduct a convergence analysis of all methods over increasing model evaluations, and finally, we conduct an in-depth performance assessment comparing accuracy and UQ of all methods.
@@
First we examine the posterior distribution resulting from all methods for $\theta=(1,10,10,10)$. Posterior distributions for $h,F,\textup{log}c$ and $b$ for the case of fewer (200) and more (1000-5400) model evaluations are shown in the left and right panels, respectively, of Figure 1 (recall that CES was run with 5400 model evaluations following methods in [11]). Due to the nonlinear and chaotic nature of the Lorenz ’96 system and correlation among parameters, the resulting posterior distributions are not expected to be Gaussian.
@@
For this particular configuration of the Lorenz ’96 system, the posterior for CES and HM are tightest around the true values, and the shape of the posterior remains very similar for either 200 (1, left panel) or 1000-5400 (1, right panel) model evaluations. For all methods, posterior distributions for parameter $b$ tend toward Gaussian, but posteriors for $F$ are more right-skewed. Despite differences in posteriors, all methods were able to capture the true parameter values (denoted by black vertical lines in Figure 1). Note also that as the number of model evaluations increases, it may not necessarily be the case that the variance of the posterior distribution will decrease. As model evaluations increase, methods such as CES and GBOED especially may explore previously under-sampled regions of parameter space, leading to a more realistic representation of uncertainty.
@@
Figure 1: Frequency histograms of posterior distributions of parameters $h,F,\log c$, and $b$ resulting from CES200, HM200, BOED200, and GBOED200 (left) and CES5400, HM1000, BOE1000, and GBOED1000 (right). Histograms represent 8092 values. True parameter values are denoted by vertical lines.
Convergence analysis
@@
Next we conducted an exercise to gain insight into convergence rates of the main methods in calibrating the Lorenz ’96 system for the same configuration ($\theta=(1,10,10,10)$). The goal of the convergence exercise was to determine how performance metrics of each method changed over increasing model evaluations. We report both score and AE every 100 evaluations up to 1000 total model evaluations for all methods. We included the intermediate (calibration) step of CES (EKS) as well for comparison.
AEs generally decreased as the number of model evaluations increased for parameter $b$, but there was considerable variability among methods for parameters $h$, $F$ and $\log c$ (Figure 2, left panel). Scores generally tended to improve with increasing model evaluations, but this was not consistent for all methods and parameters. EKS, BOED, and to some extent HM, were the only methods that demonstrated improvement in scores over increasing model evaluations (Figure 2, right panel). Scores for CES were relatively stable, while those for GBOED decreased very slightly from 100 to 1000 model evaluations.
@@
Figure 2: AEs (left) and scores (right) for CES, EKS, HM, BOED, and GBOED over increasing model evaluations for calibrating the Lorenz ’96 system using base parameters $\theta=(1,10,10,10)$ (results for parameter $c$ are calculated in log space).
While one might expect performance to improve monotonically with increasing model evaluations, convergence trends varied across parameters and methods, with performance often plateauing after an initial improvement. This behavior may reflect a combination of surrogate saturation, parameter non-identifiability, and irreducible variability associated with finite-time statistics in the chaotic Lorenz ’96 system. Notably, only EKS exhibited consistent improvement in score with increasing evaluations, which is expected given its ensemble-based contraction mechanism that directly reduces posterior variance. In contrast, methods relying on emulator-based design or filtering exhibited diminishing returns once calibration-relevant regions were adequately explored. This convergence analysis suggests that CES, GBOED/BOED, and HM may require fewer model evaluations for chaotic systems such as the Lorenz ’96 multiscale system. Finally, it is important to note that these conclusions are based on a single set of true parameters for this particular system.
Performance evaluation
@@
To obtain a fair and rigorous comparison of the main methods, we performed Bayesian inversion for $\theta_{i},i=1\dots 60$ using the Lorenz ’96 multiscale system (Eq. 10). Each component of $\theta_{i}$ was drawn from the following prior distributions based on baseline parameter values of 1, 10, log10, and 10 for $h,F$, $\log c$, and $b$, respectively:
@@
where, in this formulation, $\beta_{1}$ represents the effects of the 10 levels of $METHOD$ relative to the referencereference level, and $\beta_{2}$ represents the effects of the three levels of $PARAM$ relative to the reference level. This model can be generally expressed in matrix form as $\mathbf{Y}=\mathbf{X}\boldsymbol{\beta}+\boldsymbol{\epsilon}$, where $\mathbf{X}$ is a design matrix that encodes the levels of $METHOD$ and $PARAM$. The coefficient vector $\boldsymbol{\beta}$ contains ten coefficients for $METHOD$, three coefficients for $PARAM$, and one for the intercept. Additionally, we excluded extremely low score values (less than -50, primarily from LHS methods) from the dataset, as these could have disproportionately influenced the results. In total, 10 observations were removed. After fitting these models, we performed post-hoc pairwise comparisons of the estimated marginal means of AE by $METHOD$ to determine if differences between any two calibration methods were significantly different. Pairwise comparisons between methods were performed using Tukey’s honestly significant difference (HSD) procedure. Significance of pairwise tests was assessed at $\alpha=0.05$. All models were fitted in R using the lm() function [60]; pairwise comparisons were carried out using the emmeans [45] and modelbased [51] packages.
Results
All methods except BOED- and LHS-based approaches yielded reasonably accurate point estimates of all parameters (Figure 3A) and provided a good balance between accuracy and uncertainty quantification (Figure 3B). LHS-based methods, in particular, consistently underperformed in terms of both accuracy and UQ, emphasizing the limitations of relying solely on prior-based sampling without incorporating a data-informed calibration step (Figure 3 A and B).
Figure 3: Boxplots of AEs (A) and log score (B) for calibrated parameters $h,F,\log c$ and $b$ of the Lorenz ’96 model for all methods. Each boxplot for AE represents 60 values; values less than -50 for score were omitted. AE values were calculated using the mean of the posterior distribution. Horizontal lines within boxplots represent the median, while stars denote means.
Pairwise comparisons of estimated marginal means of AE (averaged over all parameters) from the linear AE model (Eq. 12) revealed notable performance trends. LHS-based methods (LHS200, LHS1000) and BOED200 performed significantly worse than all other methods (Table 1, Figure 4). CES5400, EKS5400, and GBOED1000 formed a top-performing group, each significantly outperforming BOED200, BOED1000, EKS200, LHS200, and LHS1000 (Figure 4, Table 1). HM methods also performed reasonably well; both HM200 and HM1000 performed similarly and achieved significantly lower AEs than BOED and LHS methods.